Lindon Company is the exclusive distributor for an automotive product that sells for $58.00 per unit and has a CM ratio of 30Outback Outfitters sells recreational equipment. One of the companys products, a small camp stove, sells for $130 per unit.Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 point? If the
Answer #1

1) variable cost per unit = 58*70% = 40.60 per unit

2) Break even unit = Fixed cost/CM per unit = 435000/17.40 = 25000 Units

Break even sales = 25000*58 = $1450000

3) Required unit = (435000+261000)/17.40 = 40000 Units

required sales = 40000*58 = $2320000

4) New break even unit = 435000/23.20 = 18750 Units

New break even sales = 18750*58 = $1087500

Required sales = (435000+261000)/.40 = $1740000